# When Geoscience Meets Foundation Models: Towards General Geoscience Artificial Intelligence System

### v3

@@ Abstract
Geoscience foundation models (GFMs) represent a revolutionary approach within Earth sciences to integrate massive cross-disciplinary data for improved simulation and understanding of Earth system dynamics. As a data-centric artificial intelligence paradigm, GFMs extract valuable insights from petabytes of both structured and unstructured data. Their versatility in task specification, diverse inputs and outputs, and multi-modal knowledge representation enable a comprehensive analysis that surpasses the capabilities of individual data sources. Critically, the scalability and generalizability of GFMs empower them to address a wide array of prediction, simulation, and decision tasks related to the intricate interactions among Earth system components. By unraveling the causal mechanisms underlying observed patterns and changes, GFMs contribute to advancing our knowledge of the Earth system and its responses to various drivers and perturbations. Collaboration between domain experts and computer scientists plays a pivotal role in fostering innovations in these invaluable tools for understanding the past, present, and future of our planet. Moreover, we introduce recent advances in GFMs including key technologies for constructing GFMs, large language models, large vision models, and large vision-language models, especially remote sensing applications. However, challenges remain in validation and verification, scalability, interpretability, knowledge representation, and addressing social bias. Going forward, the key lies in enhancing model integration, resolution, accuracy, and equity through interdisciplinary teamwork. Despite current limitations, GFMs hold great promise for providing critical insights into pressing issues including climate change, natural hazards, and sustainability through their ability to explore multiple scenarios and quantify uncertainties. Their continued evolution toward integrated, data-driven modeling holds paradigm-shifting potential for Earth science.
@@
Foundation model, Geoscience artificial intelligence, Data-centric paradigm, Causal inference, Multimodal learning
@@ I Introduction
AS Earth science enters the era of big data, artificial intelligence (AI) offers great potential not only for solving classic geoscientific problems, but also for enhancing understanding of Earth’s complex, interactive, and multiscale processes [1, 2]. The availability of massive volumes of Earth system data, already exceeding dozens of petabytes with hundreds of terabytes transmitted daily, has led to the widespread utilization of AI, including machine learning and deep learning, in data-driven Earth system science [3, 4]. For example, deep learning has been successfully applied to identify extreme weather patterns [5] to build competitive weather prediction models ranging from precipitation nowcasting [6, 7] to medium-range weather seasonal forecasts [8, 9, 10, 11], and to predict climate phenomena such as ENSO [12] or monsoon onsets [13]. Deep learning models can also automatically extract patterns and features from vast amounts of satellite and aerial imagery, which enables a wide range of applications, including environmental monitoring, land use and land cover classification, urban planning, disaster response, and climate change analysis [14, 15, 16, 17]. Moreover, generative models such as generative adversarial networks or diffusion models promise to be very valuable for bias correcting and upsampling Earth system model fields [18, 19, 20, 21]. However, fully utilizing the explosive growth and diversity of Earth system data presents two major challenges in the coming years: (i) conventional AI models have limited ability to extract knowledge from the data deluge, and (ii) supervised models cannot readily adapt to varying tasks through one training, limiting generalization. To leverage the massive amounts of Earth system data, larger models are urgently needed for data-driven Earth system science.
Foundation models—large pre-trained models representing the newest generation of AI—are trained on vast, diverse datasets and can be fine-tuned for various downstream tasks [22]. These models achieve outstanding performance across wide ranges of tasks, from answering textual queries [23] to describing images [24] or playing video games [25]. The adaptability of foundation models is a key feature that sets them apart from earlier AI models. Earlier AI models were designed for specific singular tasks and lacked the capacity to perform well on diverse tasks outside of their original purpose. In contrast, the extensive pre-training of foundation models on large and varied datasets allows them to learn rich representations of language and visual data.
Foundation models offer unprecedented capabilities, driven by the expansion of datasets, the scaling up of model size, and advancements in model architectures. For instance, as of 2023, the language model ChatGPT-4 and Claude-3 have incorporated vision, speech, and knowledge graphs, astounding users with their ability to generate text closely resembling human writing and create images and computer code from various prompts [26, 27, 23]. Moreover, many contemporary foundation models can process inputs from and produce outputs across multiple domains. Early research has explored medical foundation models capable of working with both images and text [28, 29, 30, 31]. As certain capabilities only emerge at larger model sizes, it remains difficult to predict the future accomplishments of even bigger models.
Although there have been preliminary efforts to develop geoscience foundation models (GFMs) [32, 33, 21], this shift has not yet become widespread in geoscience AI, owing to the persisting challenges in accessing expansive, diverse Earth system datasets and the complexity of the Earth system, as well as the recency of this advancement. As a result, geoscience AI models predominantly continue to be developed using a task-specific approach. For instance, a weather prediction model may be trained solely on historical climate records from a specific region, likely requiring substantial data collection efforts. This model would only be able to forecast climate change for that region over a few days, and would be unable to predict rainfall-induced shallow landslides [34]. Moreover, a remote sensing model trained on satellite images from arid regions might fail when applied to classify wetlands in Southeast Asia. The difference in vegetation, water bodies, and terrain means the model may inaccurately identify land features, leading to incorrect analyses of such ecologically sensitive areas (Aslam et al., 24; Chen et al., 2024) [35, 36]. Such narrow, task-specific models yield inflexible systems, constrained to performing predefined tasks based on the training dataset and labels. Currently, these models typically cannot adapt to new tasks or different data distributions without complete retraining. Although the task-specific approach persists as the dominant paradigm in geoscience AI, emerging foundation models may enable more flexible, multifaceted systems given sufficient accessible Earth system data.
Recent advancements in foundation models research have the potential to disrupt the prevailing task-specific paradigm in the geoscience by enabling a more versatile and adaptable approach to data analysis and prediction. Key developments include the emergence of multimodal architectures and self-supervised learning techniques that do not require explicit labels, such as language modeling and contrastive learning [37, 38]. Additionally, the advent of in-context learning capabilities allows foundation models to rapidly adapt to new tasks and data distributions. These innovations may transform how geoscientists process and interpret data, leading to enhanced accuracy and impact in critical research areas like sustainability, climate change, extreme weather events, natural hazards, and environmental monitoring [39]. As foundation models continue to evolve and increase in scale, their potential to revolutionize geoscience research and applications is immense. Adoption of these adaptable, multifaceted foundation models in lieu of narrow, static task-specific models may yield new insights and capabilities in the geoscience.
Inspired directly by foundation models outside geoscience, we identify three crucial capabilities that distinguish GFMs from conventional geoscience AI models. Firstly, adapting a GFM for new task will become much easier. These models will be able to address novel problems simply by receiving a textual explanation of the new task (flexible task specification), without requiring full retraining. Secondly, GFMs can accept varied combinations of multimodal data as input and generate diverse multimodal outputs. This enables processing of diverse data types including images, text, laboratory results, or any combination thereof. This flexible interactivity contrasts with inflexible multi-modal models that rely on predefined modalities for inputs and outputs. For example, such models must always take in a fixed set of images, text, and lab results together. Thirdly, GFMs will incorporate formal representations of geoscience system knowledge, empowering analysis of novel tasks and communication of outputs using precise professional language. Overall, GFMs are distinguished by their flexible task specification, versatile multimodal capabilities, and formal geoscience knowledge representations.
@@ II Key advantages of geoscience foundation models
Figure 1: Overview of the foundation model. a. The development of AI. Machine learning pays its attention on “how” to define the features of data and make decision using the learning algorithms, while deep learning learns features automatically by designing specific architectures. Foundation models focus on functionalities such as in-context learning emerge by using large pre-trained models. b. The characteristics of foundation models. Foundation models have five characteristics, large data, large pre-trained models, large computing force, self-training, and wide applications. c. The paradigm shift from model-centric AI to data-centric AI.
@@ II-A Foundation models: a new paradigm
Machine learning and deep learning have been widely applied for solving problems in Earth system science [3, 1, 40, 41]. As presented in Figure 1a, machine learning focuses on “how” to define the features of data and make decisions using learning algorithms, while deep learning infers “features” automatically by designing specific architectures. Foundation models focus on functionalities such as in-context learning by using large pre-trained models. They rely on massive, diverse data, large pre-trained models, and self-supervised training, yet exhibit even larger computing demands.
@@ II-A1 Data-centric AI
Drawing inspiration from foundational models beyond geoscience, we have identified three distinct features that set GFMs apart from traditional geoscience machine and deep learning models (Figure 1b). Firstly, employing GFMs for novel tasks will be as straightforward as articulating the tasks in natural language processing, such as English or other languages. GFMs will possess the capability to address previously unencountered challenges through dynamic task specification, enabling them to comprehend and tackle new tasks without the need for retraining. Secondly, GFMs can accept inputs and generate outputs utilizing various combinations of data modalities. This flexibility allows for the incorporation of different types of data, such as videos, images, sound, text, laboratory results, or any combination thereof, into the model. This flexible interactivity capability contrasts with the constraints of more stringent traditional multimodal models. Thirdly, GFMs will employ a formal representation of geoscientific knowledge, enabling it to engage in reasoning processes for previously unseen tasks and provide interpretations using geoscientifically accurate language.
@@ II-A2 Scalability and generalizability
As illustrated in Figure 1c, foundation models include large amounts of data, large models, requiring large computing capacity, self-training, all enabling wide ranges of applications, exhibiting excellent scalability and generalizability. The scalability is enabled along three key dimensions: data, model size, and computing power, and generalizability refers to the ability of a model to perform well on new tasks beyond the data it was trained on.
@@ II-B Geoscience domain knowledge
As shown in Figure 2, geoscience encompasses diverse domains spanning the physical makeup and processes of the Earth. Core geoscience domains include geology, geography, geophysics, hydrology, oceanography, meteorology, and planetary science. Each domain has accumulated extensive observations, data, and models of a wide range of Earth system components. For instance, geology involves studying the composition, structure, and history of the Earth’s crust and interior, accumulating knowledge on rock types, plate tectonics, mineral resources, and geologic time [45]. Meanwhile, meteorology focuses on the atmosphere and atmospheric phenomena like weather patterns, clouds, and precipitation. Key meteorological knowledge includes air masses, fronts, forecasting models, and climate dynamics [46]. Moreover, geography tends to explore the spatial patterns and processes of natural phenomena, such as climate change, natural disasters, and biodiversity, as well as human aspects like urbanization, cultural landscapes, and economic development. Geographers require a diverse set of knowledge, including understanding physical geography, human geography, and technical skills in geographic information systems (GIS), remote sensing, and spatial analysis [47]. Beyond domain-specific knowledge, geosciences also rely on integrative theories and principles regarding Earth systems interactions, such as the carbon cycle, the hydrologic cycle, or interactions between highly nonlinear components of the Earth system that may exhibit abrupt regime transitions [48]. Together, robust domain knowledge across the geosciences enables understanding the complex, interconnected processes that shape the Earth and its environments. Accumulating, codifying, and connecting this knowledge is an ongoing endeavor as geoscience domains continue to expand their theoretical and empirical foundations through observation, experimentation, and modeling.
@@
Figure 2: Overview of the GFM. By leveraging techniques like self-supervised learning, the GFMs undergo training using multimodal and multidimensional geoscience data models. For enhanced versatility, the data schema can be coupled with language inputs, such as text, voice, or video data. Furthermore, the GFMs require access to diverse sources of geoscientific knowledge to facilitate geoscientific inference tasks, unlocking a plethora of functionalities applicable to downstream applications. Once trained, the GFMs can dynamically perform user-specified tasks in real-time. To accomplish this, the model can retrieve contextual information from geological mapping manuscripts, knowledge graphs, databases, and other relevant sources, capitalizing on formal geoscience knowledge to deduce solutions for previously unseen tasks.
II-C Potential Applications

### v4

@@ Abstract
Artificial intelligence (AI) has significantly advanced Earth sciences, yet its full potential in to comprehensively modeling Earth’s complex dynamics remains unrealized. Geoscience foundation models (GFMs) emerge as a paradigm-shifting solution, integrating extensive cross-disciplinary data to enhance the simulation and understanding of Earth system dynamics. These data-centric AI models extract insights from petabytes of structured and unstructured data, effectively addressing the complexities of Earth systems that traditional models struggle to capture. The unique strengths of GFMs include flexible task specification, diverse input-output capabilities, and multi-modal knowledge representation, enabling analyses that surpass those of individual data sources or traditional AI methods. This review not only highlights the key advantages of GFMs, but also presents essential techniques for their construction, with a focus on transformers, pre-training, and adaptation strategies. Subsequently, we examine recent advancements in GFMs, including large language models, vision models, and vision-language models, particularly emphasizing the potential applications in remote sensing. Additionally, the review concludes with a comprehensive analysis of the challenges and future trends in GFMs, addressing five critical aspects: data integration, model complexity, uncertainty quantification, interdisciplinary collaboration, and concerns related to privacy, trust, and security. This review offers a comprehensive overview of emerging geoscientific research paradigms, emphasizing the untapped opportunities at the intersection of advanced AI techniques and geoscience. It examines major methodologies, showcases advances in large-scale models, and discusses the challenges and prospects that will shape the future landscape of GFMs. The paper highlights a dynamic field rich with possibilities, poised to unlock new insights into Earth’s complexities and further advance geoscience exploration.
@@
Artificial intelligence, geoscience foundation models (GFMs), large models, remote sensing, Earth system science.
@@ I Introduction
[…]
[lines=2]AS Earth science enters the era of big data, artificial intelligence (AI) holds significant potential not only for addressing classic geoscientific problems and enhancing understanding of Earth’s complex, interactive, and multiscale processes [1, 2]. The availability of massive volumes of Earth system data, alread exceeding dozens of petabytes with hundreds of terabytes transmitted daily, has facilitated the widespread adoption of AI technologies, in data-driven Earth system sciences [3, 4]. The utilization of AI not only enhances geological events (e.g., earthquakes and volcanic eruptions) through real-time seismic data analysis but also holds significant value in bias correction and upsampling of Earth system models [5, 1, 6, 7, 8]. Machine learning approaches support the exploration and mapping of natural resources (e.g., minerals and oil), thereby increasing accuracy and minimizing the ecological impacts of mining and drilling operations [9, 10, 11]. Moreover, the application of deep learning algorithms in climate modeling and weather forecasting plays a crucial role in understanding the effects of climate change, refining weather predictions, and enhancing disaster management and preparedness [12, 13, 14, 15, 16, 17]. Additionally, the advanced capabilities of AI in processing and interpreting extensive amounts of satellite imagery and sensor data also makes a substantial contribution to environmental monitoring and conservation, which enables a diverse array of critical applications, including environmental monitoring, land cover classification, urban planning, disaster response, and climate change analysis [18, 19, 20, 21].
The successful applications of AI technologies in earth science present formidable challenges and inherent limitations that necessitate careful consideration. A predominant issue is the dependency on extensive high-quality datasets, which are frequently scarce or incomplete in geoscience disciplines stemming from the complex and variable nature of Earth systems [22, 23]. Furthermore, the generalization capability of AI models is frequently compromised when transitioning between different geological contexts. This dependence on specific training data can result in significant inaccuracies when models are applied to novel or poorly understood environments [24, 25, 26]. Another critical limitation is the interpretability of AI outputs. For AI-driven insights to be effectively incorporated into geoscience research and practical applications, they must be transparent and comprehensible, thereby fostering trust and adoption among geoscientists [27, 28, 29]. To address these challenges, the development of larger, more robust models is imperative for advancing data-driven Earth system science.
Recently, foundation models have emerged as pivotal tools for tackling many complex data-centric tasks through pre-training on vast datasets and subsequent fine-tuning for specific applications [30]. Foundation models exhibit unparalleled emergent capabilities, including contextual reasoning, complex problem-solving, and zero-shot adaptability across diverse tasks, which are driven by advanced algorithms, extensive training on vast datasets, and sophisticated neural network architectures [31, 32, 33, 34]. The generalization capability of foundation models distinctly sets them apart from traditional AI models, which were confined to specific, singular tasks and exhibited limited performance on diverse tasks beyond their original scope. In contrast, foundation models undergo extensive pre-training on expansive, varied datasets, enabling them to develop comprehensive representations of language and visual data [35]. For instance, as of 2024, the large language models (LLMs), such as ChatGPT-4 [36] and Claude-3/3.5 [37, 38], have integrated capabilities in vision, speech, and knowledge graphs, impressing users with their ability to produce text indistinguishable from human writing and generate images and computer code from diverse prompts [16, 39, 33, 40]. Moreover, many contemporary foundation models are capable of processing both images and text, finding applications in diverse fields such as medicine, transportation, remote sensing, and urban planning [41, 42, 43, 44, 45, 46]. Given the transformative potential of such technologies, it is imperative to develop specialized foundation models for geoscience, which can revolutionize the field by providing deep, actionable insights into complex Earth systems.
Although there have been preliminary efforts to develop geoscience foundation models (GFMs) [47, 48, 8, 49, 45, 50, 46], widespread adoption within geoscience AI remains limited. The slow progression in GFMs is attributed to persistent challenges in accessing and processing multi-source and multi-modal geoscience datasets, the inherent complexity of Earth systems, and the inherent limitations of current AI technologies [51, 52, 53]. Although the task-specific approach persists as the dominant paradigm in geoscience AI. These innovations in GFMs may transform how geoscientists process and interpret data, leading to enhanced accuracy and impact in critical research areas like sustainability, climate change, extreme weather events, natural hazards, and environmental monitoring [54, 55, 56]. Therefore, the adoption of adaptable, multifaceted foundation models in place of narrow, static task-specific models has the potential to not only provide new insights and capabilities to Earth science, but also to transform our understanding of complex Earth systems, drive innovation in interdisciplinary research, and address global challenges with greater precision and foresight.
This study methodically traces the evolution of AI applications within the earth sciences, highlighting the limitations of traditional AI methodologies. It underscores the growing importance of foundation models in geosciences, analyzes the transformative impact of these models on geoscience research paradigms, and explores potential research directions for GFMs. The study also elaborates on the unique capabilities of GFMs, such as their adaptability to diverse tasks, their proficiency in processing multimodal data, and their seamless integration of complex geoscientific knowledge. Together, these characteristics demonstrate the potential of GFMs to redefine task specification, enhance multimodal capabilities, and strengthen geoscientific knowledge representations, thereby establishing a new paradigm and standard for geoscience AI.
I-A Existing Surveys and Reviews
Table I provides an overview of existing surveys and reviews related to foundation models for geoscience. These works have explored various aspects of foundation models in geoscience, including remote sensing, urban science, vision-language models, generative artificial intelligence, geophysical foundation models, and geospatial location embedding. These works are published in 2023 and 2024. For example, Jiao et al. [57] contributed to the field with their proposed brain-inspired framework for remote sensing foundation models. In a Chinese-language publication, Zhang’s team [58] offered a comprehensive review of pre-training techniques for large remote sensing models. The landscape of smart cities is enriched by the work’s [55] introduction of Urban Foundation Models. The authors in [45] delved into the intersection of vision-language models and remote sensing in their examination. The potential of generative AI and large language models in geoscience is explored in depth by the research group led by Hadid et al. [59]. A thought-provoking discussion on geophysical foundation models in exploration geophysics was presented in the perspective piece by the work [60]. Zhou et al. spearheaded a thorough review [61] that sheds light on vision-language geo-foundation models. The authors in [62] break new ground with their investigation into geospatial location embedding in large language models. However, there is still a need for a comprehensive survey that integrates these diverse perspectives and provides a holistic view of the current state of the field. This paper aims to fill this gap by providing a comprehensive review of the key advantages, techniques, and advances in GFMs. We also discuss the challenges and future trends of GFMs in geoscience, highlighting the potential of these models to revolutionize Earth system science.
TABLE I: Existing Surveys and Reviews Related to Foundation Models for Geoscience
Year Publication Summary Advs. Techs. Adv.
2023 Jiao et al. [57] This survey examines foundation models in remote sensing, comparing their performance across various tasks and datasets, and proposes a brain-inspired framework for remote sensing foundation models while identifying open research problems.
2023 Zhang et al. [58] This review examines pre-training techniques for large remote sensing models, analyzing supervised and unsupervised single-modal models as well as visual-text joint multimodal models, and discusses future prospects for integrating domain knowledge, enhancing generalization, expanding applications, and reducing data costs. (in Chinese)
2024 Zhang et al. [55] This paper introduces urban foundation models (UFMs), proposes a data-centric taxonomy for current UFM-related works, presents a framework for their future development, and explores their potential applications in smart cities.
2024 Li et al. [45] This review comprehensively examines the application of vision-language models (VLMs) in remote sensing, covering various tasks such as image captioning, text-based image generation and retrieval, visual question answering, and scene analysis, while discussing current progress, challenges, and future research directions in this emerging field.
2024 Hadid et al. [59] This survey explores the applications and potential of generative artificial intelligence (GAI) and large language models in geoscience, discussing various GAI models used in the field, their applications, challenges, and future promises in addressing Earth system dynamics and related scientific domains.
2024 Liu et al. [60] This perspective discusses the potential applications, research directions, and challenges of developing geophysical foundation models (GeoFMs) in exploration geophysics and related sub-disciplines, proposing strategies for their construction and exploring new research paradigms in the field.
2024 Zhou et al. [61] This paper provides a comprehensive review of vision-language geo-foundation models (VLGFMs), covering their background, core technologies, applications, and future research directions in the field of multimodal geospatial data analysis.
2024 Tucker et al. [62] This review explores geospatial location embedding (GLE) in large language models, identifying four main themes and proposing a spatial foundation/language mdel (SLM) framework to advance spatial artificial intelligence systems (SPAIS) with embedded spatial knowledge.
2024 This work A comprehensive review of geoscience foundation models (GFMs) from the view of key advantages, key techniques, and key advances.
I-B Contributions and Organization
Though there exists a growing body of literature on foundation models in geoscience, the field is still in its infancy, with many open research questions and challenges. Table I provides an overview of existing surveys and reviews related to foundation models for geoscience. These works have explored various aspects of foundation models in geoscience, including remote sensing, urban science, vision-language models, generative artificial intelligence, geophysical foundation models, and geospatial location embedding. However, there is still a need for a comprehensive survey that integrates these diverse perspectives and provides a holistic view of the current state of the field. This paper aims to fill this gap by providing a comprehensive review of the key advantages, techniques, and advances in geoscience foundation models. We also discuss the challenges and future trends of GFMs in geoscience, highlighting the potential of these models to revolutionize Earth system science. The main contributions of this paper are as follows
  • This review provides a comprehensive overview of geoscience foundation models (GFMs) from the perspective of key advantages, key techniques, and key advances. The study also discusses the challenges and future trends of GFMs in geoscience, emphasizing the potential of these models to revolutionize Earth system science.
  • This study introduces a new paradigm by integrating AI foundation models with specialized geoscience knowledge. It addresses the unique challenges of geoscience data, including its spatial-temporal complexity and interdisciplinary nature. The research aims to develop a general geoscience AI system capable of handling diverse tasks across space, air, ground, and ocean domains. This integration enhances the scalability and generalizability of AI models in geoscience, potentially transforming how researchers approach complex earth science problems.
  • We present a comprehensive review of the key advances in GFMs, including large language models (LLMs), large vision models (LVMs), and large vision-language models (LVLMs), with a particular emphasis on their applications in remote sensing. We also discuss the potential of GFMs to revolutionize Earth system science by providing deep, actionable insights into complex Earth systems.
  • This study contributes a thorough analysis of challenges and future trends in GFMs. We identify key obstacles including data complexity, model interpretability, uncertainty quantification, interdisciplinary collaboration, and privacy, trust and security concerns. This forward-looking perspective guides strategic development and innovation in the rapidly evolving field of geoscience AI.
Figure 1: Organization of the review.
The remainder of this paper is organized as follows. Section II discusses the key advantages of GFMs, including their data-centric paradigm, scalability, generalizability, dynamic interactions, and causal inference. Section III introduces the geoscience domain knowledge and the potential applications of GFMs in Earth science. Section IV presents the recent advances in GFMs, including large language models, large vision models, and large vision-language models, especially remote sensing applications. Section V discusses the challenges and future trends of GFMs. Finally, Section VI concludes the paper.
@@ II Key Advantages of Geoscience Foundation Models
In this section, we discuss the key advantages of geoscience foundation models (GFMs) from the view of a new paradigm, Geoscience domain knowledge, and application domains. Data-centric AI, ccalability and generalizability, and dynamic interactions make GFMs a new paradigm compared to traditional AI. Geoscience domain knowledge including publiocations, geological data, and knowledge graph further enhance the capabilities of GFMs. The application domain of GFMs in Earth science are also discussed, including space, air, ground, and ocean domains.
@@ II-A Foundation Models: A New Paradigm
Machine learning and deep learning have been widely applied for solving problems in Earth system science [3, 1, 63, 64]. Machine learning focuses on “how” to define the features of data and make decisions using learning algorithms, while deep learning infers “features” automatically by designing specific architectures. Foundation models focus on functionalities such as in-context learning by using large pre-trained models. They rely on massive, diverse data, large pre-trained models, and self-supervised training, yet exhibit even larger computing demands.
@@ II-A1 Data-Centric AI
Drawing inspiration from foundational models beyond geoscience, we have identified three distinct features that set GFMs apart from traditional geoscience machine and deep learning models. Firstly, employing GFMs for novel tasks will be as straightforward as articulating the tasks in natural language processing, such as English or other languages. GFMs will possess the capability to address previously unencountered challenges through dynamic task specification, enabling them to comprehend and tackle new tasks without the need for retraining. Secondly, GFMs can accept inputs and generate outputs utilizing various combinations of data modalities. This flexibility allows for the incorporation of different types of data, such as videos, images, sound, text, laboratory results, or any combination thereof, into the model. This flexible interactivity capability contrasts with the constraints of more stringent traditional multimodal models. Thirdly, GFMs will employ a formal representation of geoscientific knowledge, enabling it to engage in reasoning processes for previously unseen tasks and provide interpretations using geoscientifically accurate language.
@@ II-A2 Scalability and Generalizability
Foundation models include large amounts of data, large models, requiring large computing capacity, self-training, all enabling wide ranges of applications, exhibiting excellent scalability and generalizability. The scalability is enabled along three key dimensions: data, model size, and computing power, and generalizability refers to the ability of a model to perform well on new tasks beyond the data it was trained on.
@@ II-B Geoscience Domain Knowledge
As shown in Fig. 2, geoscience encompasses diverse domains spanning the physical makeup and processes of the Earth. Core geoscience domains include geology, geography, geophysics, hydrology, oceanography, meteorology, and planetary science. Each domain has accumulated extensive observations, data, and models of a wide range of Earth system components. For instance, geology involves studying the composition, structure, and history of the Earth’s crust and interior, accumulating knowledge on rock types, plate tectonics, mineral resources, and geologic time [68]. Meanwhile, meteorology focuses on the atmosphere and atmospheric phenomena like weather patterns, clouds, and precipitation. Key meteorological knowledge includes air masses, fronts, forecasting models, and climate dynamics [69]. Moreover, geography tends to explore the spatial patterns and processes of natural phenomena, such as climate change, natural disasters, and biodiversity, as well as human aspects like urbanization, cultural landscapes, and economic development. Geographers require a diverse set of knowledge, including understanding physical geography, human geography, and technical skills in geographic information systems (GIS), remote sensing, and spatial analysis [70].
Beyond domain-specific knowledge, geosciences also rely on integrative theories and principles regarding Earth systems interactions, such as the carbon cycle, the hydrologic cycle, or interactions between highly nonlinear components of the Earth system that may exhibit abrupt regime transitions [71]. Together, robust domain knowledge across the geosciences enables understanding the complex, interconnected processes that shape the Earth and its environments. Accumulating, codifying, and connecting this knowledge is an ongoing endeavor as geoscience domains continue to expand their theoretical and empirical foundations through observation, experimentation, and modeling.
@@
Figure 2: Overview of the Geoscience foundation model (GFM). By leveraging techniques like self-supervised learning, the GFMs undergo training using multimodal and multidimensional geoscience data models. For enhanced versatility, the data schema can be coupled with language inputs, such as text, voice, or video data. Furthermore, the GFMs require access to diverse sources of geoscientific knowledge to facilitate geoscientific inference tasks, unlocking a plethora of functionalities applicable to downstream applications. Once trained, the GFMs can dynamically perform user-specified tasks in real time. To accomplish this, the model can retrieve contextual information from geological mapping manuscripts, knowledge graphs, databases, and other relevant sources, capitalizing on formal geoscience knowledge to deduce solutions for previously unseen tasks.
II-C Application Domain